# Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star forming galaxies at z5.2 with the N2CLS

### v1

@@ Appendix D Notes about individual sources and JWST cutouts
Below, we provide a detailed description of each individual source identified at z5.2 (but HDF850.1, GN10 and N2GN_1_13). The JWST photometry measurements follow a similar approach to that described by Weibel et al. (2024), with some modifications as outlined in Section 2.8 of B25. In Figure 7, we present the multi-wavelength JWST postage stamps, where available.
N2GN_1_20 is detected by SMA. SMA position lies at 1.2 arcsec from S-CANDELS J123550.35+621041.8. It has a Ks band flux of 0.75 Jy, resulting in a very red Ks to 3.6 m color, , interpreted as a deep break at 4000 Å combined to a large extinction, that leads to a (Liu et al. 2018). Using New-Hyperz, we obtained =5.16, the broad error bars being the result of a degenerated solution in the parameter space between redshit, extinction and age of the stellar population. This source is not covered by JWST.
@@
N2GN_1_23: is detected by JWST/FRESCO at =5.179 and is the so-called S3 source extensively discussed in Xiao et al. (2024). We have a firm identification of our N2CLS source through our NOEMA follow-up program (see B25) with the NOEMA counterpart lying at a distance of 0.167 arcsec. The NIKA2 source has both a SCUBA2 450 and 850 m counterparts but the NOEMA position lies at 3.6 arcsec from the 450 m source Number 78 of Barger et al. (2022), identified with a galaxy at z=0.29 (see their Table 2). The 850 m source is identified with SMM123656621205 (Number 80), which is at 2.8 arcsec from the NOEMA position, and with a redshift estimate of 3.75 (Cowie et al. 2017). The high-z source is clearly blended with a nearby object and flux measurements from Herschel and SCUBA2 450 m cannot be used to construct the multi-wavelength SED. The NOEMA source has both a 1.4 GHz (Owen 2018) and 10 GHz (Jiménez-Andrade et al. 2024) radio counterpart.
@@
N2GN_1_26 is identified with a high- HST galaxy thanks to SMA data. Kodra et al. (2023) give a photometric redshift . The FIR SED is well defined by the PACS, SCUBA2, SMA, and NIKA2 data. The 24 m flux is well matched by an intense 3.3 m PAH feature. There is a 10 GHz source at the HST position. This source is newly observed and detected by the PANORAMIC survey with F277W, F356W. Using New-Hyperz, it has a photometric redshift of 4.87, consistent with previous findings at 1 level.
N2GN_1_33: corresponds to the SMA GOODS-N source number 13 (Cowie et al. 2017). Three sources lie inside the SMA beam: a bright optical galaxy, a very faint optical galaxy revealed by JWST, and a radio source which lies in the middle of the two optical sources. The bright and faint optical galaxies are respectively at 1.21 and 0.56 arcsec from the SMA position. We chose the closest one as the identification. This is confirmed by the 10 GHz low- and high-resolution maps which have a clear source at this position. In B25, it has ; using New-Hyperz, we obtain , in perfect agreement with previous findings.
@@
N2GN_1_43 is identified with NOEMA. It is a K-band dark galaxy. It is detected by JWST at in the F277W band, in F200W, and tentatively by HST at shorter wavelengths. Although not in the IRAC catalog, it is clearly seen in the maps at 4.5, 5.8 and 8 m. It has also a clear 10 GHz signal (Jiménez-Andrade et al. 2024) at the JWST source position. A fit with CIGALE in B25 places it at with a 4000 Å break between the two JWST measurements mentioned here. The photometric redshift from New-Hyperz is 5.31, consistent with previous estimates, with a degenerated solution in the parameter space between redshift and extinction leading to broad error bars.
@@
N2GN_1_57: The VLA source, coinciding with GN4 26261 (MIOP program) at z=0.9, is likely not the right counterpart as it lies at 4.9 arcsec from the N2CLS position. Closer to the N2CLS position (2.2 arcsec), the source HRG14 J123731.66+621616.7 (Finkelstein et al. 2015) (also detected by the SHARDS middle-band survey) is at a photometric redshift of z=5.20.06 in Arrabal Haro et al. (2018) and z=5.30 in Kodra et al. (2023). It is detected in the four IRAC channels and has a pronounced 4000 Å break at m. The photometric redshift obtained with New-Hyperz is z=5.18, consistent with previous findings.
@@
N2GN_1_61 is not only optically-dark but also IRAC-dark ([4.5]¿24.6 mag at 5). The accurate position is given by NOEMA (see B25) revealing a single source with a flux of 69884Jy at 1.18 mm, in perfect agreement with the NIKA2 flux. JWST grism observations from FRESCO detected one emission line, consistent with H at (M. Xiao, priv. comm.). It is observed by JWST with F090W, F115W, F182M, F210M, F356W, F444W filters, but only detected by F356W and F444W (see Fig. 7). The observed SED consists of only three data points (NIKA2 1.2 mm, SCUBA2 850 m and JWST F444W), but the availability of the redshift allows us to perform a fit. Derived parameters have large uncertainties (but are consistent between MAGPHYS-highz and CIGALE).

### v2

@@ Appendix C Notes about individual sources and JWST cutouts
Below, we provide a detailed description of each individual source identified at z5.2 in the overdensity (but HDF850.1, GN10, and N2GN_1_13). The JWST photometry measurements follow a similar approach to that described by weibel2024, with some modifications as outlined in Sect. 2.8 of B25. The multi-wavelength cutouts of all sources are shown in B25. We complement those cutouts with the JWST postage stamps, where available, in Fig. 7.
@@
N2GN_1_23: is detected by JWST FRESCO at =5.179 and is the so-called S3 source extensively discussed in Xiao2024. We have a firm identification of our N2CLS source through our NOEMA follow-up programme (see B25) with the NOEMA counterpart lying at a distance of 0.167 arcsec. The NIKA2 source has both a SCUBA2 450 and a 850 m counterpart but the NOEMA position lies at 3.6 arcsec from the 450 m source Number 78 of Barger2022, identified with a galaxy at z=0.29 (see their Table 2). The 850 m source is identified with SMM123656621205 (Number 80), which is at 2.8 arcsec from the NOEMA position, and with a redshift estimate of 3.75 (Cowie2017). The high-z source is clearly blended with a nearby object and flux measurements from Herschel and SCUBA2 450 m cannot be used to construct the multi-wavelength SED. The NOEMA source has both a 1.4 GHz (owen2018) and 10 GHz (jimenez2024) radio counterpart.
@@
N2GN_1_26 is identified with a high- HST galaxy thanks to SMA data. kodra2023 give a photometric redshift . The far-IR SED is well defined by the PACS, SCUBA2, SMA, and NIKA2 data. The 24 m flux is well matched by an intense 3.3 m polycyclic aromatic hydrocarbon feature. There is a 10 GHz source at the HST position. This source is newly observed and detected by the PANORAMIC survey with F277W, F356W. Using New-Hyperz, it has a photometric redshift of 4.87, consistent with previous findings at 1 level.
@@
N2GN_1_43 is identified with NOEMA. It is a K-band dark galaxy. It is detected by JWST at in the F277W band, in F200W, and tentatively by HST at shorter wavelengths. Although not in the IRAC catalogue, it is clearly seen in the maps at 4.5, 5.8 and 8 m. It has also a clear 10 GHz signal (jimenez2024) at the JWST source position. A fit with CIGALE in B25 places it at with a 4000 Å break between the two JWST measurements mentioned here. The photometric redshift from New-Hyperz is 5.31, consistent with previous estimates, with a degenerated solution in the parameter space between redshift and extinction leading to broad error bars.
@@
N2GN_1_57: The VLA source, coinciding with GN4 26261 (MIOP programme) at z=0.9, is likely not the right counterpart as it lies at 4.9 arcsec from the N2CLS position. Closer to the N2CLS position (2.2 arcsec), the source HRG14 J123731.66+621616.7 (also detected by the SHARDS middle-band survey; finkelstein2015) is at a photometric redshift of z=5.20.06 in arrabal_haro2018 and z=5.30 in kodra2023. It is detected in the four IRAC channels and has a pronounced 4000 Å break at m. The photometric redshift obtained with New-Hyperz is z=5.18, consistent with previous findings.
@@
N2GN_1_61 is not only optically dark but also IRAC-dark ([4.5]¿24.6 mag at 5). The accurate position is given by NOEMA (see B25) revealing a single source with a flux of 69884Jy at 1.18 mm, in perfect agreement with the NIKA2 flux. JWST grism observations from FRESCO detected one emission line, consistent with H at (M. Xiao, priv. comm.). It is observed by JWST with F090W, F115W, F182M, F210M, F356W, F444W filters, but only detected by F356W and F444W (see Fig. 7). The observed SED consists of only three data points (NIKA2 1.2 mm, SCUBA2 850 m and JWST F444W), but the availability of the redshift allows us to perform a fit. Derived parameters have large uncertainties (but are consistent between MAGPHYS-highz and CIGALE).
Appendix D Another N2CLS candidate at z5.2
We identified an additional N2CLS galaxy at z5.2 that lies close to the overdense structure. N2GN_1_20 is detected by SMA. SMA position (RA=12:35:50.22 ; Dec=62:10:42.4) lies at 1.2 arcsec from S-CANDELS J123550.35+621041.8, which is an optically dark galaxy. It has a Ks band flux of 0.75 Jy, resulting in a very red 3.6 m to Ks colour (), interpreted as a deep break at 4000 Å combined to a large extinction, that leads to a (liu2018). Using New-Hyperz, we obtained =5.16, the broad error bars being the result of a degenerated solution in the parameter space between redshift, extinction and age of the stellar population. This source is not covered by JWST.
We show in Fig. 8 the best fit from CIGALE, assuming z=5.33. The parameters derived from the fit are: SFR = 2160611 M⊙ yr-1; AV = 2.70.3; M⋆ = (30366)1010 M⊙; and Mdust = (23.319.70) M⊙. Combining the derived SFR and M⋆, and applying the tacconi2020 scaling relations, we obtain Mgas= 1031010 M⊙ and = 476 Myr. Assuming its photometric redshift, the galaxy is extremely massive. As such, it is lying between the popesso2023 and speagle2014 parametrisations of the MS.
It is also lying above the model expectations for the GOODS-N survey volume (Fig. 3), requiring =0.7 for the full sky area. This discrepancy could arise from assumptions related to the IMF. For example, a top-heavy IMF could yield lower stellar masses by dex, resulting in slightly reduced values of . It could also arise from an overestimation of its photometric redshifts. A substantially lower true redshift compared to the photometric estimate may result in a lower best-fit stellar mass, thereby reducing the required value of . To test this hypothesis, we modelled the SED of N2GN_1_20 with CIGALE assuming z=3 (see Fig. 8). We obtain a stellar mass of (3.150.34) M⊙, essentially comparable to that inferred at z=5.33. It is important to note that the fit at z=3 is not merely a rescaling of the same template at a different redshift: CIGALE recomputes the fit and searches for an optimal solution at the imposed z. The reduced is 30% higher at z=3 than at z=5.33. Such a high stellar mass at z=3 remains above model expectations for the GOODS-N field area, and would still require =0.12 for the full sky. If confirmed at z=5.33, the extremely high mass could reflect the presence of an unusually dense concentration of massive halos in this region.
We also compare the SED of N2GN_1_20 with those of the eight galaxies discussed in the main text, using different normalisations (Fig. 9). We do not find any indication that N2GN_1_20 deviates significantly from the others.
As shown in Fig. 5, N2GN_1_20 lies outside the core of the overdensity. Its distance to GN10 is RA=11.8 cMpc, Dec=-8.1 cMpc, and LOSdistance= 14.9 cMpc. Further investigation is required to determine whether this optically dark galaxy truly lies at high redshifts, to understand its unusually high stellar mass and efficiency, and to assess whether it may trace an extended region of the overdensity.